A random sample of 160 commercial customers of Charter Communication revealed that 32 had paid their bills within a month of billing. The corresponding 90% confidence interval for the true proportion of the customers who paid their bills within a month is? a. 0.153 to 0.247. b. 0.138 to 0.262. c. 0.144 to 0.256. d. 0.148 to 0.252.
Answer #1

Answer- d. 0.148 to 0.252.

Reason-

Sample proportion formula=

The interval is p ± z[p(1 - p)/n]^1/2 =

Where p= 20%=0.20, n=162, z value in 90% confidence interval = 1.64

=.20 + (1.64)[(.20)(.80)/160]^1/2= 0.250

and .20 - (1.64)[(.20)(.80)/160]^1/2= 0.148